CH_{3}^{+}: A NEW TOOL FOR STUDYING THE ENIGMA OF INTERSTELLAR CH^{+}?

\underline{BENJAMIN J. McCALL}, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801; THOMAS R. GEBALLE, Gemini Observatory, Hilo, HI 96720; TAKESHI OKA, Department of Astronomy & Astrophysics and Department of Chemistry, University of Chicago, Chicago, IL 60637.

Although CH^{+} was one of the very first interstellar molecules to be identified (in the early 1940s), its ubiquitous presence in the diffuse interstellar medium remains an enigma. CH^{+} is thought to be rapidly destroyed by the abundant species H, H_{2}, and e^{−}, but there is no known efficient formation mechanism at low temperatures. In steady state, one would therefore expect a very low abundance of CH^{+}, but yet strong CH^{+} lines are seen in nearly all diffuse cloud sightlines. Various solutions to this paradox have been proposed, and most of them invoke shocks (especially magnetohydrodynamic shocks) to drive the endothermic reaction C^{+} + H_{2} → CH^{+} + H.

Regardless of the mechanism by which it is produced, CH^{+} will undergo fast hydrogen abstraction reactions to form CH_{2}^{+} and then CH_{3}^{+}, if it is in an environment with a large fraction of the hydrogen in molecular form. Since CH_{3}^{+} reacts only very slowly with H_{2}, by radiative association to form CH_{5}^{+}, it is destroyed primarily by dissociative recombination with electrons. A simple steady state analysis suggests that CH_{3}^{+} may be over 10 times more abundant than CH^{+}, although this factor depends on the molecular fraction and the electron temperature.

In order to provide observational constraints on the environment in which CH^{+} is observed, we have searched for the infrared ν_{3} band of CH_{3}^{+} at 3.2 μm in diffuse cloud sources. We will present our upper limits on the equivalent width of the CH_{3}^{+} lines, discuss the inferred limits on the CH_{3}^{+} column density, and offer some speculations about the conditions in the regions where CH^{+} exists.